A quantity of gasoline has a mass of 3450 kg and volume of 5m^3, calculate its density.
1
Expert's answer
2020-04-22T10:03:33-0400

The density of a gasoline

"\rho=\frac{m}{V}=\frac{3450\:\rm kg}{5\:\rm m^3}=690\:\rm kg/m^3."